# FeDaL: Federated Dataset Learning for Time Series Foundation Models

### v1

@@ 4.2 Downstream Task Adaption as Time Series Foundation Models
To evaluate the generality of FeDaL-pretrained TSFMs, we test on diverse downstream tasks: long-/short-term forecasting, imputation, classification, and anomaly detection. Pretraining is conducted on LOTSA dataset woo2024unifiedtraininguniversaltime  (231B time points across 174 client-specific subsets spanning nine domains). We use 1000 rounds with 10 local steps, and SGD with learning rate $1e^{-2}$. Other hyperparameters follow Section˜4.1, and data preprocessing details are provided in Appendix˜B, full results in Appendix˜C.
@@
Table 2: Long-term forecasting results (averaged across horizons $\{96,192,336,720\}$ for ETT-series and Weather, and $\{24,36,48,60\}$ for ILI) in full-shot and few-shot settings (5% and 10% training data). Bold: best; Underline: second best. None of these datasets were included in pretraining. * denotes a “former“ suffix.
@@
Models FeDaL (Ours) FFTS FedAvg TimeMixer Time-LLM GPT4TS PatchTST TimesNet DLinear Fed.* Auto.* Stationary LightTS In.* Re.*
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Full-shot Long-term Forecasting
ETTh1 0.380 0.409 0.391 0.412 0.412 0.431 0.448 0.443 0.408 0.423 0.465 0.455 0.413 0.430 0.458 0.45 0.422 0.437 0.440 0.460 0.496 0.487 0.570 0.537 0.491 0.479 1.040 0.795 1.029 0.805
ETTh2 0.334 0.377 0.334 0.389 0.340 0.382 0.364 0.394 0.334 0.383 0.381 0.412 0.330 0.379 0.414 0.427 0.431 0.446 0.437 0.449 0.450 0.459 0.526 0.516 0.602 0.543 4.431 1.729 6.736 2.191
ETTm1 0.319 0.365 0.323 0.374 0.333 0.367 0.381 0.395 0.329 0.372 0.388 0.403 0.351 0.380 0.400 0.406 0.357 0.378 0.448 0.452 0.588 0.517 0.481 0.456 0.435 0.437 0.961 0.734 0.799 0.671
ETTm2 0.261 0.319 0.253 0.314 0.254 0.316 0.275 0.323 0.251 0.313 0.284 0.339 0.255 0.315 0.291 0.333 0.267 0.333 0.305 0.439 0.327 0.371 0.306 0.347 0.409 0.436 1.410 0.810 1.479 0.915
Weather 0.213 0.255 0.217 0.256 0.226 0.260 0.241 0.272 0.225 0.257 0.237 0.270 0.255 0.264 0.259 0.287 0.248 0.300 0.309 0.360 0.338 0.382 0.288 0.314 0.261 0.312 0.634 0.548 0.803 0.656
ILI 1.355 0.773 1.389 0.798 1.410 0.800 2.039 0.899 1.435 0.801 1.925 0.903 1.443 0.797 2.139 0.931 2.169 1.041 2.847 1.144 3.006 1.161 2.077 0.914 7.382 2.003 5.137 1.544 4.724 1.445
Few-shot Long-term Forecasting (5%)
ETTh1 0.592 0.511 0.601 0.520 0.630 0.537 1.193 0.740 0.627 0.543 0.681 0.560 0.694 0.569 0.925 0.647 0.750 0.611 0.658 0.562 0.722 0.598 0.943 0.646 1.451 0.903 1.225 0.817 1.241 0.835
ETTh2 0.350 0.391 0.360 0.400 0.388 0.416 0.502 0.493 0.382 0.418 0.400 0.433 0.827 0.615 0.439 0.448 0.694 0.577 0.463 0.454 0.441 0.457 0.470 0.489 3.206 1.268 3.922 1.653 3.527 1.472
ETTm1 0.400 0.401 0.417 0.415 0.415 0.414 0.566 0.503 0.425 0.434 0.472 0.450 0.526 0.476 0.717 0.561 0.400 0.417 0.730 0.592 0.796 0.620 0.857 0.598 1.123 0.765 1.163 0.791 1.264 0.826
ETTm2 0.240 0.307 0.253 0.299 0.258 0.308 0.340 0.371 0.274 0.323 0.308 0.346 0.314 0.352 0.344 0.372 0.399 0.426 0.381 0.404 0.381 0.404 0.341 0.372 1.415 0.871 3.658 1.489 3.581 1.487
Weather 0.222 0.286 0.241 0.295 0.242 0.295 0.260 0.303 0.260 0.309 0.263 0.301 0.269 0.303 0.298 0.318 0.263 0.308 0.309 0.353 0.309 0.353 0.327 0.328 0.305 0.345 0.584 0.527 0.447 0.453
Few-shot Long-term Forecasting (10%)
ETTh1 0.540 0.505 0.549 0.513 0.581 0.520 0.613 0.520 0.556 0.522 0.590 0.525 0.633 0.542 0.869 0.628 0.691 0.600 0.639 0.561 0.702 0.596 0.915 0.639 1.375 0.877 1.199 0.808 1.249 0.833
ETTh2 0.352 0.388 0.360 0.391 0.382 0.400 0.402 0.433 0.370 0.394 0.397 0.421 0.415 0.431 0.479 0.465 0.605 0.538 0.466 0.475 0.488 0.499 0.462 0.455 2.655 1.160 3.872 1.513 3.485 1.486
ETTm1 0.389 0.407 0.389 0.410 0.410 0.428 0.487 0.461 0.404 0.427 0.464 0.441 0.501 0.466 0.677 0.537 0.411 0.429 0.722 0.605 0.802 0.628 0.797 0.578 0.971 0.705 1.192 0.821 1.426 0.856
ETTm2 0.261 0.312 0.264 0.315 0.289 0.339 0.311 0.367 0.277 0.323 0.293 0.335 0.296 0.343 0.320 0.353 0.316 0.368 0.463 0.488 1.342 0.930 0.332 0.366 0.987 0.756 3.370 1.440 3.978 1.587
Weather 0.207 0.256 0.214 0.261 0.240 0.278 0.242 0.281 0.234 0.273 0.238 0.275 0.242 0.279 0.279 0.301 0.241 0.283 0.284 0.324 0.300 0.342 0.318 0.323 0.289 0.322 0.597 0.495 0.546 0.469
Average 0.396 0.390 0.405 0.396 0.422 0.404 0.578 0.435 0.419 0.406 0.478 0.427 0.484 0.434 0.577 0.469 0.534 0.469 0.607 0.505 0.708 0.513 0.620 0.491 1.562 0.766 2.142 1.010 2.359 1.059
$1^{st}$ Count 30 2 0 0 2 0 1 0 1 0 0 0 0 0 0
Long-Term Forecasting
@@
Table 3: Average zero-shot forecasting performance across horizons $\{96,192,336,720\}$ for observation lengths $\{512,1024,2048,3072\}$. Bold: best; Underline: second best. Full results in Table˜14, Appendix˜C.
Metrics FeDaL (Ours) FFTS FedAvg Moiraismall Moiraibase Moirailarge TimesFM Moment Chronossmall Chronosbase Chronoslarge
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.407 0.429 0.425 0.437 0.438 0.449 0.428 0.427 0.417 0.419 0.480 0.439 0.473 0.443 0.683 0.566 0.545 0.472 0.591 0.468 0.588 0.466
ETTh2 0.361 0.382 0.370 0.374 0.390 0.401 0.361 0.384 0.362 0.382 0.367 0.377 0.392 0.406 0.361 0.409 0.424 0.430 0.405 0.410 0.455 0.427
ETTm1 0.360 0.390 0.364 0.420 0.378 0.410 0.436 0.410 0.406 0.385 0.422 0.391 0.433 0.418 0.670 0.536 0.640 0.499 0.645 0.500 0.555 0.465
ETTm2 0.292 0.341 0.317 0.362 0.322 0.365 0.307 0.347 0.311 0.337 0.329 0.343 0.328 0.346 0.316 0.365 0.349 0.380 0.310 0.350 0.295 0.338
Weather 0.255 0.284 0.262 0.300 0.277 0.305 0.275 0.286 0.287 0.281 0.264 0.273 - - 0.294 0.326 0.300 0.318 0.292 0.315 0.279 0.306
Average 0.335 0.365 0.348 0.378 0.361 0.386 0.361 0.371 0.357 0.361 0.372 0.373 0.407 0.428 0.402 0.440 0.452 0.420 0.457 0.409 0.414 0.396
1st Count 6 1 0 1 4 1 0 1 0 1 2
Short-Term Forecasting
To evaluate the effectiveness of the learned TSFM via our FeDaL in short-term forecasting tasks, we conduct experiments on the M4 dataset, following the protocols of zhou2023one . As shown in Table 4, FeDaL significantly outperforms baselines. While it performs slightly below FFTS on SMAPE by a narrow margin of 0.07%, FeDaL reduces MASE by 2% to 38% and achieves 5% to 22% improvements in SMAPE and OWA. These further confirm the ability of FeDaL to learn cross-domain representations that generalize effectively in short-term forecasting tasks.
Table 4: Average short-term forecasting results on M4 dataset. Bold: best; Underline: second best. None of these datasets were included in pretraining. * denotes a “former“ suffix. Full results are provide in Table˜16.
Models
@@
Imputation
Imputation aims to reconstruct missing values in time series based on partially observed data. We evaluate our method on five widely used real-world datasets, including four ETT datasets, and Weather, where the data-missing is common. Following the evaluation protocol of GPT4TS zhou2023one , we randomly mask time points at varying ratios to simulate different levels of corruption. As shown in Table˜5, the TSFM trained via our FeDaL consistently outperforms all baselines across datasets. Compared to the state-of-the-art centralized model GPT4TS, FeDaL reduces the average MSE by 22.84%. When compared to FL baselines FFTS and FedAvg, FeDaL achieves 12.64% and 27.62% relative MSE reductions, respectively. These demonstrate the effectiveness of FeDaL in learning common temporal patterns that generalize well even under missing-data conditions.
Table 5: Average imputation performance for randomly masked time series (input length 96), evaluated across four mask ratios $\{12.5\%,25\%,37.5\%,50\%\}$. Bold: best; Underline: second best. Full results in Table˜18.
Models FeDaL FFTS FedAvg TimeMixer GPT4TS PatchTST TimesNet DLinear Fed. Auto. Stationary LightTS In.
@@
ETTh1 0.022 0.090 0.024 0.090 0.034 0.120 0.036 0.123 0.028 0.105 0.047 0.14 0.12 0.253 0.027 0.107 0.104 0.218 0.093 0.206 0.062 0.177 0.051 0.15 0.071 0.188
ETTh2 0.018 0.071 0.017 0.074 0.026 0.098 0.028 0.102 0.021 0.084 0.029 0.102 0.208 0.327 0.022 0.088 0.046 0.151 0.096 0.208 0.101 0.215 0.029 0.105 0.156 0.292
ETTm1 0.054 0.147 0.058 0.160 0.054 0.154 0.073 0.192 0.069 0.173 0.115 0.224 0.202 0.329 0.078 0.187 0.284 0.373 0.201 0.306 0.117 0.246 0.103 0.214 0.161 0.279
ETTm2 0.034 0.106 0.046 0.135 0.062 0.154 0.038 0.120 0.048 0.141 0.065 0.163 0.367 0.436 0.049 0.146 0.119 0.25 0.142 0.259 0.163 0.279 0.055 0.156 0.337 0.452
Weather 0.024 0.048 0.029 0.059 0.034 0.050 0.039 0.083 0.031 0.056 0.03 0.054 0.076 0.171 0.03 0.054 0.055 0.117 0.052 0.11 0.099 0.203 0.031 0.057 0.045 0.104
Average 0.030 0.092 0.035 0.104 0.042 0.115 0.043 0.124 0.039 0.112 0.057 0.137 0.195 0.303 0.041 0.116 0.122 0.222 0.117 0.218 0.108 0.224 0.054 0.136 0.154 0.263
1st Count 11 2 0 0 0 0 0 0 0 0 0 0 0
Anomaly Detection
Anomaly detection focuses on identifying abnormal patterns or segments. We evaluate FeDaL on five widely used datasets, including SMD, MSL, SMAP, SwaT, and PSM, following the evaluation protocols as GPT4TS zhou2023one to ensure fair comparison. The results as shown in Table˜6, the FeDaL-trained TSFM achieves the best overall performance across all datasets. It consistently outperforms prior state-of-the-art models, including ModernTCN and the general-purpose TSFM Moment, by 2.40% and 5.17%, respectively. Notably, from an FL perspective, FeDaL also delivers significant gains over FL baselines such as FedAvg and FFTS by 0.96% and 3.69%, respectively. These demonstrate FeDaL’s strong capability to capture global temporal invariances and its superior generalization to complex anomaly detection across heterogeneous client distributions.
Table 6: Anomaly detection results. We calculate the F1-score (%) for each dataset and statics the average F1-score. Bold: the best, Underline: the second best. * denotes a “former“ suffix. Full results are in Table˜19.
@@
Avg. F1
88.70
87.86
85.54
84.34
86.72
86.62
86.34
84.97
84.23
82.46
82.08
84.26
82.57
80.50
78.83
77.31
Classification
To assess cross-task generalization, we evaluate the learned TSFM on time series classification using 10 UEA bagnall2018uea and 91 UCR dau2019ucr subsets spanning diverse domains. We adopt Linear Probing method by attaching a linear classifier to the frozen TSFM, directly measuring the quality of learned representations. As shown in Figure˜5, our FeDaL-trained TSFM consistently outperforms all baselines, including task-specific models, GPT4TS (fine-tuned), and Moment (generalized). Notably, FeDaL surpasses FL baselines such as FedAvg and FFTS by 6.4%/6.1% on UEA and 9.5%/3.2% on UCR, highlighting its ability to learn domain-invariant features and mitigate cross-client biases.
Figure 5: Classification results. Best viewed in color and with zoom.

### v2

@@ 4.2 Adaption as General Time Series Foundation Models
To evaluate the generality of FeDaL-pretrained TSFMs, we test on diverse downstream tasks including long-/short-term deterministic forecasting, imputation, classification, and anomaly detection. Pretraining is conducted on LOTSA (Woo et al., 2024), which consists of 231B time points partitioned into 174 dataset-specific clients, naturally aligning with the federated setting. For non–zero-shot task (expect long-term forecasting), we fine-tune the trained model for only one epoch. Data preprocessing details are provided in Appendix B, and full results are included in Appendix C.
@@
Table 3: Full-shot Long-term forecasting results (averaged across horizons $\{96,192,336,720\}$ for ETT-series and Weather, and $\{24,36,48,60\}$ for ILI). Bold: best; Underline: second best.
@@
Models FeDaL FFTS FedAvg TimeMixer Time-LLM GPT4TS PatchTST TimesNet DLinear Fedformer Autoformer Stationary LightTS
ETTh1 0.380 0.409 0.391 0.412 0.412 0.431 0.448 0.443 0.408 0.423 0.465 0.455 0.413 0.430 0.458 0.450 0.422 0.437 0.440 0.460 0.496 0.487 0.570 0.537 0.491 0.479
ETTh2 0.334 0.377 0.334 0.389 0.340 0.382 0.364 0.394 0.334 0.383 0.381 0.412 0.330 0.379 0.414 0.427 0.431 0.446 0.437 0.449 0.450 0.459 0.526 0.516 0.602 0.543
ETTm1 0.319 0.365 0.323 0.374 0.333 0.367 0.381 0.395 0.329 0.372 0.388 0.403 0.351 0.380 0.400 0.406 0.357 0.378 0.448 0.452 0.588 0.517 0.481 0.456 0.435 0.437
ETTm2 0.261 0.319 0.253 0.314 0.254 0.316 0.275 0.323 0.251 0.313 0.284 0.339 0.255 0.315 0.291 0.333 0.267 0.333 0.305 0.439 0.327 0.371 0.306 0.347 0.409 0.436
Weather 0.213 0.255 0.217 0.256 0.226 0.260 0.241 0.272 0.225 0.257 0.237 0.270 0.255 0.264 0.259 0.287 0.248 0.300 0.309 0.360 0.338 0.382 0.288 0.314 0.261 0.312
ILI 1.355 0.773 1.389 0.798 1.410 0.800 2.039 0.899 1.435 0.801 1.925 0.903 1.443 0.797 2.139 0.931 2.169 1.041 2.847 1.144 3.006 1.161 2.077 0.914 7.382 2.003
1st Count 9 0 0 0 2 0 1 0 0 1 2 2 2
Table 4: Average zero-shot forecasting performance across horizons $\{96,192,336,720\}$ for observation lengths $\{512,1024,2048,3072\}$. Bold: best; Underline: second best. Full results in Table 19, Appendix C.
Models FL Pretraining Centralized Time Series Foundation Models
Metrics FeDaL (Ours) FFTS FedAvg Moiraib Moirail TimesFM Moment Chronosb Chronosl Time-MoEb Time-MoEl Time-MoEu
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.407 0.429 0.425 0.437 0.438 0.427 0.417 0.419 0.480 0.439 0.473 0.443 0.683 0.566 0.591 0.468 0.588 0.466 0.400 0.424 0.394 0.419 0.412 0.426
ETTh2 0.361 0.382 0.370 0.374 0.390 0.401 0.362 0.382 0.367 0.377 0.392 0.406 0.361 0.409 0.405 0.410 0.455 0.427 0.366 0.404 0.405 0.415 0.371 0.399
ETTm1 0.360 0.390 0.364 0.420 0.378 0.410 0.406 0.385 0.422 0.391 0.433 0.418 0.670 0.536 0.645 0.500 0.555 0.465 0.394 0.415 0.376 0.405 0.356 0.391
ETTm2 0.292 0.341 0.317 0.362 0.322 0.365 0.311 0.337 0.329 0.343 0.328 0.346 0.316 0.365 0.310 0.350 0.295 0.338 0.317 0.365 0.316 0.361 0.288 0.344
Weather 0.255 0.284 0.262 0.300 0.277 0.305 0.287 0.281 0.264 0.273 0.294 0.326 0.292 0.315 0.279 0.306 0.265 0.297 0.270 0.300 0.256 0.288
Avg. 0.335 0.365 0.348 0.379 0.361 0.382 0.357 0.361 0.372 0.365 0.407 0.403 0.465 0.440 0.449 0.409 0.434 0.400 0.348 0.381 0.352 0.380 0.337 0.370
1st Count 3 1 0 4 1 0 1 0 1 0 0 2
$2^{nd}$ Count 4 0 0 2 2 0 0 0 1 2 2 2
Data Scale 231B 231B 231B 231B 231B 100B 1.13B 84B 84B 300B 300B 300B
Params# 28.42M 31.18M 28.42M 84M 91M 200M 385M 200M 710M 113M 453M 2.4B
Long-Term Forecasting.
@@
Short-Term Forecasting.
To evaluate the effectiveness of FeDaL-trained TSFM in short-term forecasting tasks, we conduct experiments on M4 dataset, following the protocols of GPT4TS (Zhou et al., 2023). As shown in Table 5, FeDaL significantly outperforms baselines. While it performs slightly below FFTS on SMAPE by a narrow margin of 0.07%, FeDaL reduces MASE by 2% to 38% and achieves 5% to 22% improvements in SMAPE and OWA. These further confirm the ability of FeDaL to learn cross-dataset representations that generalize effectively in forecasting.
Table 5: Average short-term forecasting results on M4 dataset. Bold: best; Underline: second best. None of these datasets were included in pretraining. * denotes a “former“ suffix. Full results are provide in Table 21.
Model
@@
Imputation.
Imputation evaluates a model’s ability to reconstruct missing values. We conduct experiments on five widely used datasets, including four ETT and Weather, where missingness is common. Following FFTS (Chen et al., 2025), we simulate varying corruption levels by randomly masking points. As shown in Table 6, FeDaL consistently outperforms all baselines. Compared to the strongest FL baselines, it reduces MSE by 12.64% over FFTS and 27.62% over FedAvg. Against centralized TSFMs, FeDaL achieves 22.84% lower error than GPT4TS and also surpasses the general-purpose TSFM Moment, despite Moment being trained on a much larger centralized corpus. These confirm that FeDaL not only leverages distributed data effectively but also achieves competitive or superior generalization to centralized TSFMs under missing-data conditions.
Table 6: Average imputation performance for randomly masked time series, evaluated across four mask ratios $\{12.5\%,25\%,37.5\%,50\%\}$. Bold: best; Underline: second best. Full results in Table 23.
Model FeDaL FFTS FedAvg Moment TimeMixer GPT4TS PatchTST TimesNet DLinear Fedformer Autoformer NS LightTS
@@
ETTh1 0.022 0.090 0.024 0.090 0.034 0.120 0.024 0.094 0.036 0.123 0.028 0.105 0.047 0.140 0.120 0.253 0.027 0.107 0.104 0.218 0.093 0.206 0.062 0.177 0.051 0.150
ETTh2 0.018 0.071 0.017 0.074 0.026 0.098 0.021 0.080 0.028 0.102 0.021 0.084 0.029 0.102 0.208 0.327 0.022 0.088 0.046 0.151 0.096 0.208 0.101 0.215 0.029 0.105
ETTm1 0.054 0.147 0.058 0.160 0.054 0.154 0.062 0.167 0.073 0.192 0.069 0.173 0.115 0.224 0.202 0.329 0.078 0.187 0.284 0.373 0.201 0.306 0.117 0.246 0.103 0.214
ETTm2 0.034 0.106 0.046 0.135 0.062 0.154 0.042 0.130 0.038 0.120 0.048 0.141 0.065 0.163 0.367 0.436 0.049 0.146 0.119 0.250 0.142 0.259 0.163 0.279 0.055 0.156
Weather 0.024 0.048 0.029 0.059 0.034 0.050 0.029 0.061 0.039 0.083 0.031 0.056 0.030 0.054 0.076 0.171 0.030 0.054 0.055 0.117 0.052 0.110 0.099 0.203 0.031 0.057
1st Count 9 2 1 0 0 0 0 0 0 0 0 0 0
Anomaly Detection.
We evaluate FeDaL-trained TSFM on five widely used multivariate datasets (SMD, MSL, SMAP, SwaT, PSM), following Moment (Goswami et al., 2024) protocols for fair comparison. As shown in Table 7, FeDaL achieves the best overall performance across all datasets. It surpasses strong centralized baselines such as ModernTCN and TSFM Moment by 2.40% and 5.17%, respectively, and further outperforms FL methods including FedAvg and FFTS by 0.96% and 3.69%. These results highlight our FeDaL‘s effectiveness in capturing global temporal invariance and its superior generalization to complex anomaly detection tasks.
Table 7: Anomaly detection results. We calculate the F1-score (%) for each dataset and statics the average F1-score. Bold: the best, Underline: the second best. * denotes a “former“ suffix. Full results are in Table 24.
@@
1st Count
3
0
0
0
1
0
0
0
0
0
0
0
0
0
0
0
Figure 5: Classification results. Full results in Table 25 & Table 26. Best viewed in color and with zoom.
Classification.
Following the standard evaluation protocol from (Zhang et al., 2025a), We evaluate the FeDaL-trained TSFM on time series classification using 10 UEA (Bagnall et al., 2018) and 91 UCR (Dau et al., 2019) subsets spanning diverse domains. We adopt Linear Probing-based adaption method by attaching a linear classifier to the frozen TSFM, directly measuring the quality of learned representations. As shown in Figure 5, our FeDaL-trained TSFM consistently outperforms all baselines, including task-specific models, GPT4TS (fine-tuned), and Moment. Notably, FeDaL surpasses FL baselines such as FedAvg and FFTS by 6.4%/6.1% on UEA and 9.5%/3.2% on UCR, highlighting its ability to learn domain-invariant features and mitigate cross-client biases.
Figure 6: T-SNE visualization of Core-set.
Core-Set Analysis.
To better understand the effect of core-set tuning, we visualize the representation spaces of three key stages using t-SNE in Figure 6. The initial core-sets cluster closely around the original client data, indicating that they effectively approximate local distributions, but may also retain sensitive fine-grained patterns. After applying Fourier perturbation and semantic alignment, the refined core-sets move away from the raw data distribution while remaining semantically aligned. This suggests that our perturbation strategy can partially obfuscate fine-grained details and reduce direct data leakage risks, while preserving high-level temporal semantics that are useful for global alignment. As a result, the uploaded core-sets serve as compact, privacy-aware summaries that enable effective server-side tuning without requiring access to raw client data. More detailed analysis of the t-SNE visualizations for our proposed core-set tuning is provided in Figure 8.